# Enabling Large-Scale and High-Precision Fluid Simulations on Near-Term Quantum Computers

### v1

@@ 6 Conclusion and Outlook
In this paper, we propose a novel approach to Computational Fluid Dynamics (CFD) leveraging the capabilities of near-term quantum computers, along with several enhancements to existing methods to enable effective CFD simulations. First, we introduce a high-precision quantum linear solver suitable for Noisy Intermediate-Scale Quantum (NISQ) devices, called Iterative-QLS. This solver effectively mitigates errors that typically arise in quantum computing processes. Our simulation of a 2D Poiseuille flow on a near-term quantum computer demonstrates a relative error of less than $0.2\%$, even lower than the gate error of the quantum computer. To address the challenges in quantum state tomography, we develop a sparse tomography algorithm, which removes the dependency on Quantum Random Access Memory (QRAM) inherent in $l_{\infty}$-norm tomography. Second, we introduce the subspace method to scale the Iterative-QLS method for larger CFD problems, enhancing solution convergence and resulting in a quantum-classical hybrid approach. By employing the subspace method, we successfully simulate unsteady acoustic wave propagation on a $41\times 41$ grid using a NISQ computer, marking a significant advancement toward scalable quantum-enhanced CFD solutions.
@@ C.2 Convergence analysis
Therefore $j_{0}$ has the order of $j_{0}\in\widetilde{\mathcal{O}}(\sqrt{b})$ to allow the truncation with a bounded error. Finally, we see that there must exist a polynomial $p(x)=\sum_{i=0}^{j_{0}}a_{i}\mathcal{T}_{i}(x)$ with at most $i$-th order where $j_{0}\in\mathcal{O}(\kappa\log 1/\epsilon)$. With $\boldsymbol{x}=A^{-1}\boldsymbol{b}$ and $\tilde{\boldsymbol{x}}=p(A)\boldsymbol{b}$, we can conclude that $\lVert\hat{\boldsymbol{x}}-\boldsymbol{x}\rVert\leq\epsilon\lVert\boldsymbol{x}\rVert$.
@@
One of the most common cases for applying the subspace method is to use a constant subspace’s order $m$ and the restarting scheme. It keeps updating the residual with a more precise solution and finally obtains the solution. From this idea, we show the convergence of the Sub-VQLS with two steps: first, the convergence can be satisfied with an existing solution from the subspace $\mathcal{K}_{m}$; second, the convergence can be satisfied with a fixed residual produced by VQLS method.
@@
The Sub-VQLS method produces a solution with $\epsilon$ error as long as QLS finds a solution better than $\boldsymbol{0}$ as the initial guess. That is, we can solve the linear system with precision $\max(\mathcal{O}(e^{-m^{2}/\kappa^{2}}),\epsilon_{q})<1$ such that $C=\langle\psi|A^{\dagger}(I-|b\rangle\langle b|)A|\psi\rangle\leq\epsilon_{q}$. The number of iterations is bounded by
@@
It is obvious to see when $\kappa$ is larger, we are required to have a better $\epsilon_{q}$ (or extra iterations to the subspace method). We should note that this is only a lower bound for the convergence of the VQLS method. That is, by satisfying $\epsilon_{q}<1$, one can always guarantee the convergence of the solution. In other words, Thm. 5 provides an insight into our method that the VQLS does not always require to compute an exact solution. Also, in the realistic test, the convergence can be faster.
@@ Appendix D Theory details of sparse tomography
In this section, we will provide a rigorous proof of the sparse tomography given in the main text. Firstly, we show the definition of quantum sparse sampling.
@@
For a $N$-dimension real-valued quantum state $|x\rangle=\sum_{i=0}^{N-1}x_{i}|i\rangle$, we produce a classical vector $\hat{\boldsymbol{x}}$ that satisfies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{\infty}<\epsilon$ with a number of copies of the quantum state. The classical vector has $k$ nonzero entries such that $k\ll N$. Here, real-valued means that every $x_{i}$ is a real number.
@@
Given access of unitary $U$ such that $U|0\rangle=|x\rangle$ and its controlled version, there exists an algorithm that produces a $l_{\infty}$ sampling for a real-valued quantum state with $\mathcal{O}(\log d/\epsilon^{2})$ query complexity. The algorithm requires a constant number of queries to QRAM and an extra qubit.
@@
There exists an algorithm that produces a sparse sampling for a $N$-dimensional real-valued quantum state with $\epsilon$ error with $\mathcal{O}(\log N/\epsilon^{2})$ time complexity. The algorithm can run without using QRAM and any extra qubit, i.e. all procedures are only based on measurements on different bases.
@@
For the number of measurements $M$, and defining $\delta=1/\sqrt{36\log N/\epsilon^{2}}$, we use the Hoeffding inequality and obtain
@@
Given that each edge is successfully measured with a constant success rate $p$. we can traverse this graph randomly in $v$ steps and determine all signs successfully with a probability of $p^{v}$. Repeating this random traversal $k$ times will result in a success rate of approximately $1-\left(1-p^{v}\right)^{k}\sim 1-\left(v\left(1-p\right)\right)^{k}$. Therefore, we can choose $k=(\log\delta)/\left(\log{v\left(1-p\right)}\right)$ to ensure the algorithm succeeds with a probability of $1-\delta$, with the number of measurements being $\mathcal{O}(Mvk)$, where $M$ still represents the number of measurements for each Hamiltonian, and $\mathcal{O}(vk)$ denotes the number of Hamiltonians that need to be computed.
While $p$ is generated from the Hoeffding inequality with a number of measurements $M$, where $p\geq 1-2\exp(-M\delta^{2}/2)$, we can derive an explicit sample complexity for this algorithm, given by
@@ Appendix E Additional experiments on the scalability of Iterative-QLS
The performance of Iterative-QLS to solve 2D incompressible Poiseuille flow running on a noise-free quantum circuit simulator is studied firstly, which indicates the scalability under ideal conditions. Fig. 10(a) and (b) respectively show the variation of residuals and loss values. It can be seen that iterative VQLS can successfully solve the system with 2 qubits but fails when using more than 2 qubits, even though a noticeable decrease in loss values can still be observed.

### v2

@@ 5 Conclusion and Outlook
In this paper, we propose a novel approach to Computational Fluid Dynamics (CFD) leveraging the capabilities of near-term quantum computers, along with several enhancements to existing methods to enable effective CFD simulations. First, we introduce a high-precision quantum linear solver suitable for Noisy Intermediate-Scale Quantum (NISQ) devices, called Iterative-QLS. This solver effectively mitigates errors that typically arise in quantum computing processes. Our simulation of a 2D Poiseuille flow on a near-term quantum computer demonstrates a relative error of less than $0.2\%$, even lower than the gate error of the quantum computer. To address the challenges in quantum state tomography, we develop a sparse tomography algorithm, which removes the dependency on quantum random access memory inherent in $l_{\infty}$-norm tomography. Second, we introduce the subspace method to scale the Iterative-QLS method for larger CFD problems, enhancing solution convergence and resulting in a quantum-classical hybrid approach. By employing the subspace method, we successfully simulate unsteady acoustic wave propagation on a $41\times 41$ grid using a NISQ computer, marking a significant advancement toward scalable quantum-enhanced CFD solutions.
@@ C.2 Convergence analysis
Therefore, $j_{0}$ has the order of $j_{0}\in\widetilde{\mathcal{O}}(\sqrt{b})$ to allow the truncation with a bounded error. Finally, we see that there must exist a polynomial $p(x)=\sum_{i=0}^{j_{0}}a_{i}\mathcal{T}_{i}(x)$ with at most $i$-th order where $j_{0}\in\mathcal{O}(\kappa\log 1/\epsilon)$. With $\boldsymbol{x}=A^{-1}\boldsymbol{b}$ and $\tilde{\boldsymbol{x}}=p(A)\boldsymbol{b}$, we can conclude that $\lVert\hat{\boldsymbol{x}}-\boldsymbol{x}\rVert\leq\epsilon\lVert\boldsymbol{x}\rVert$.
@@
One of the most common cases for applying the subspace method is to use a constant subspace’s order $m$ and the restarting scheme. It keeps updating the residual with a more precise solution and finally obtains the solution. From this idea, we show the convergence of the subspace method with two steps: first, the convergence can be satisfied with an existing solution from the subspace $\mathcal{K}_{m}$; second, the convergence can be satisfied with a fixed residual produced by VQLS method.
@@
The sbuspace method produces a solution with $\epsilon$ error as long as QLS finds a solution better than $\boldsymbol{0}$ as the initial guess. That is, we can solve the linear system with precision $\max(\mathcal{O}(e^{-m^{2}/\kappa^{2}}),\epsilon_{q})<1$ such that $C=\langle\psi|A^{\dagger}(I-|b\rangle\langle b|)A|\psi\rangle\leq\epsilon_{q}$. The number of iterations is bounded by
@@
It is obvious to see when $\kappa$ is larger, we are required to have a better $\epsilon_{q}$ (or extra iterations to the subspace method). We should note that this is only a lower bound for the convergence of the QLS. That is, by satisfying $\epsilon_{q}<1$, one can always guarantee the convergence of the solution. In other words, Thm. 5 provides an insight into our method that the QLS is not always required to compute an exact solution. Also, in the realistic test, the convergence can be faster.
@@ Appendix D Theory details of sparse tomography
In this section, we will provide rigorous proof of the sparse tomography given in the main text. Firstly, we show the definition of quantum sparse sampling.
@@
For an $N$-dimension real-valued quantum state $|x\rangle=\sum_{i=0}^{N-1}x_{i}|i\rangle$, we produce a classical vector $\hat{\boldsymbol{x}}$ that satisfies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{\infty}<\epsilon$ with a number of copies of the quantum state. The classical vector has $k$ nonzero entries such that $k\ll N$. Here, real-valued means that every $x_{i}$ is a real number.
@@
Given access of unitary $U$ such that $U|0\rangle=|x\rangle$ and its controlled version, there exists an algorithm that produces an $l_{\infty}$ sampling for a real-valued quantum state with $\mathcal{O}(\log d/\epsilon^{2})$ query complexity. The algorithm requires a constant number of queries to QRAM and an extra qubit.
@@
There exists an algorithm that produces a sparse sampling for an $N$-dimensional real-valued quantum state with $\epsilon$ error with $\mathcal{O}(\log N/\epsilon^{2})$ time complexity. The algorithm can run without using QRAM and any extra qubit, i.e. all procedures are only based on measurements on different bases.
@@
For the number of measurements $M$, and defining $\delta=1/\sqrt{36\log N/\epsilon^{2}}$, we use the Hoeffding’s inequality and obtain
@@
Given that each edge is successfully measured with a constant success rate $p$, we can traverse this graph randomly in $v$ steps and determine all signs successfully with a probability of $p^{v}$. Repeating this random traversal $k$ times will result in a success rate of approximately $1-\left(1-p^{v}\right)^{k}\sim 1-\left(v\left(1-p\right)\right)^{k}$. Therefore, we can choose $k=(\log\delta)/\left(\log{v\left(1-p\right)}\right)$ to ensure the algorithm succeeds with a probability of $1-\delta$, with the number of measurements being $\mathcal{O}(Mvk)$, where $M$ still represents the number of measurements for each Hamiltonian, and $\mathcal{O}(vk)$ denotes the number of Hamiltonians that need to be computed.
While $p$ is generated from Hoeffding’s inequality with a number of measurements $M$, where $p\geq 1-2\exp(-M\delta^{2}/2)$, we can derive an explicit sample complexity for this algorithm, given by
@@ Appendix E Additional experiments on the scalability of Iterative-QLS
The performance of Iterative-QLS to solve 2D incompressible Poiseuille flow running on a noise-free quantum circuit simulator is studied first, which indicates the scalability under ideal conditions. Fig. 8(a) and (b) respectively show the variation of residuals and loss values. It can be seen that iterative VQLS can successfully solve the system with 2 qubits but fails when using more than 2 qubits, even though a noticeable decrease in loss values can still be observed.
@@
Appendix F Examples for the discussion about scalability
Here, we demonstrate two numerical examples to show how our algorithm can be enhanced by a better quantum computer. Firstly, we conduct quantum circuit simulations for Poiseuille flow on a classical computer with different qubit errors. The qubit error is defined as the strength of the depolarizing error channel appended after every quantum gate. In Fig. 10, we can see that the count of the circuit execution significantly decreases with a decreasing error rate. Therefore, our method promises a faster execution with a better qubit quality that is utilized in the experiment.
Secondly, we argue that the integration of the subspace method on the flow simulation is promising when more qubits are provided to enable a larger subspace size $m=2^{n}$ where $n$ is the qubit and $m$ is the size of the subspace. In Fig. 11, we solve a matrix of $1024$ dimension with $m$ varies from $2^{4}$ to $2^{10}$. We plot the final residual and the number of steps of subspace, which corresponds to the predicted circuit execution time on a quantum computer. The result implies the problem can be divided into two regimes. At $n\leq 8$, where the subspace size is far less than the dimension, we call it the “NISQ regime”, representing the case when we have a very limited number of high-quality qubits. At $n\geq 9$, where the subspace size approaches the matrix size, we call it the “FTQC regime”, representing the case when we are able to put a full flow field on a quantum computer.
In the NISQ regime, we find the number of subspace steps follows a nearly exponential dependence on the number of qubits, while the system converges to a similar residual. This represents that the investment of every additional qubit will result in exponential computing power. In the FTQC regime, the subspace size approaches the dimension of the original matrix. As a result, the residual decreases significantly because it almost becomes a direct inversion algorithm. The number of subspace steps also decreases significantly when it enters such a regime. This represents that even in the FTQC regime, it is meaningful to explore the potential of integration of the subspace method and quantum computing before direct inversion is allowed.
Refer to caption NISQ Regime FTQC Regime
Figure 11: Results of solving a 1024-dimension matrix with different subspace sizes. (a) The residual of the final solution after convergence. (b) The number of subspace steps.